(1 point) NOTE: When using interval notation in WebWork, remember that: You use INF for oo and -INF for -0. And use U f
Answer #1

Page No. Date. f(x) = -1.5x²-sx-2 f(20) = - 3X-5 10 = 23-5/3 tl(x)2-3 to 7 KO fal + fulani Hence fis concave up on uy جای I